OpenStudy (anonymous):

Use the functions m(x) = 5x + 4 and n(x) = 6x − 9 to complete the function operations listed below. Part A: Find (m + n)(x). Show your work. (3 points) Part B: Find (m ⋅ n)(x). Show your work. (3 points) Part C: Find m[n(x)]. Show your work. (4 points)

OpenStudy (anonymous):

okay so basically what part A is saying is find m(x) + n(x) part B is saying m(x) * n(x) part C is saying plug n(x) equation for x in the m(x) equation

OpenStudy (anonymous):

so for part A do i just use the number for m that does not have a x in it?

OpenStudy (anonymous):

so basically its saying (5x+4) + (6x-9)

OpenStudy (jdoe0001):

\(\large { (m+n)(x)\implies m(x)+n(x) \\ \quad \\ (m\cdot n)(x)\implies m(x)\cdot n(x) \\ \quad \\ m[n(x)]\implies \begin{cases} m(x) = 5x + 4\\{\color{brown}{ n(x) }} = 6x - 9 \end{cases}\qquad m[n(x)]=5[{\color{brown}{ n(x) }}]+4 }\)

OpenStudy (anonymous):

idk if that answers ur quesiton

OpenStudy (anonymous):

i got 11x-5?

OpenStudy (anonymous):

@ilikemathandcoding

OpenStudy (anonymous):

omg no way so did I :D

OpenStudy (anonymous):

so what do i put for part A?

OpenStudy (anonymous):

the answer which u asked me about

OpenStudy (anonymous):

ok now part b

OpenStudy (anonymous):

tell me what you get read over my first comment and try it out

OpenStudy (anonymous):

so would i just do the same thing but multiply them instead of adding?